Zeronknight (zeronknight):

Determine the Greatest Common Factor and simplify the expression. 24x + 32x2 - 26x A. 2x(12+16x -13) B. 3x(8 + 9x - 13) C. 8x(3x + 4x2 - 3x) D. 4x(6 + 8x - 4)

Zeronknight (zeronknight):

@imqwerty

OpenStudy (jtug6):

I can help if you'd like :p

Zeronknight (zeronknight):

Yeah, could you explain this?

OpenStudy (jtug6):

Sure! So, first thing is that its asking for the GREATEST common factor of the expression 24x + 32x^2 - 26x. What common and biggest term can go into 24, 32, and 26?

Zeronknight (zeronknight):

Hmm ok. 4?

OpenStudy (jtug6):

mmm not quite. 24/4 = 6 so that checks, 32/4 = 8 so that checks, but 26/4 = 6.5 and we're looking for whole numbers. Perhaps something a bit smaller than 4? ;p

Zeronknight (zeronknight):

2 goes into all things even.

Zeronknight (zeronknight):

24/2 = 12 32 / 2 = 16 26 / 2 = 13

OpenStudy (jtug6):

Yes! 4 cant work but 2 fits into all of them. What else is common with that expression, numbers aside?

Zeronknight (zeronknight):

I'm not so sure what else is, or what you mean by that. You mean a variable that appears the most?

OpenStudy (jtug6):

Yeah. Isn't there an x-variable in all of the expression?

Zeronknight (zeronknight):

Yes, so the GCF is 2 and X

OpenStudy (jtug6):

yes sir. so, pull that out of the expression and what happens to the inside?

Zeronknight (zeronknight):

Not so sure...:\

Zeronknight (zeronknight):

Are you say line up the problem with 2 and X?

OpenStudy (jtug6):

Basically we found the GCF which is 2x so we can yank it out of 24x + 32x^2 -26x so that it looks like 2x(___ + ____ - ____)

OpenStudy (jtug6):

but whats in the underscores? :P

Zeronknight (zeronknight):

The numbers.

Zeronknight (zeronknight):

The other numbers that aren't 2x

Zeronknight (zeronknight):

But what about the numbers with X in them?

OpenStudy (jtug6):

yeah. youre essentially diving 2x out of all three since you pulled it out

OpenStudy (jtug6):

well, for the inside we have to divide by 2x so it becomes (24x)/(2x) + (32x^2)/(2x) - 26x/2x.

Zeronknight (zeronknight):

So from all of this, we find the greatest GCF out of all of them, and we take it out. We divide the GCF by the number stated above in the problem 26 / 2 = 13 etc. So the answer would be A?

OpenStudy (jtug6):

Yeah. The reason you would want to divide is because by pulling that 2x out you're now multiplying it by the inside. if we left the inside alone and we multiplied the 2x through the unchanged inside the expression we would get wouldnt match the original one.

OpenStudy (jtug6):

so 2x(24x + 32x^2 - 26x) would be equal to 48x^2 + 64x^3 - 52x^2 which is not equal to the original expression 2x(24x/2x + 32x^2/2x - 26x/2x) satisfies that we pulled the 2x out

Zeronknight (zeronknight):

Ohhh

OpenStudy (jtug6):

yup

OpenStudy (jtug6):

Just curious is this Algebra I or II?